OpenStudy (anonymous):

I am so desperate, please help! i have to factor this equation. 21r^2 - rs - 2s^2 As well as solve this one in any way: x^4-21x^2-100=0 Please! The questions are due in 10-15 minutes!

OpenStudy (anonymous):

(3r-s) (7r+2s) for the first one. (x-5) (x+5) (x^2+4) = 0 x = 5 and x = -5 for the second one. :)

OpenStudy (anonymous):

tlau, how did you get the answer?

OpenStudy (anonymous):

Factor by inspection. Valid method.

jhonyy9 (jhonyy9):

x^4 -21x^2 -100 =0 note x^2=y so than y^2 -21y -100=0 y_1,_2=(21+/- sqrt(441+400))/2=(21+/- sqrt(841))/2 =

OpenStudy (anonymous):

wouldn't the answer to the second one also be -4 as well?

OpenStudy (anonymous):

You can't square root a negative number. Have you done imaginary numbers?

OpenStudy (anonymous):

x^2 = -4 No real number x can be squared to get -4. Any number squared positive or negative is positive.

OpenStudy (anonymous):

I did 2 years ago. i'm having to relearn everything as I am starting pre-cal, but now I understand. Thank you very much!

OpenStudy (anonymous):

No problem. :D